### v2

# Taking quantisation seriously: a farewell to waves

Abstract
The dual wave-particle nature of quantum objects is a notoriously unintuitive feature of quantum theories. However, it is often deemed essential, due to otherwise discrete particles exhibiting diffraction and interference. Troublingly, Landé and Lévy-Leblond have shown that de Broglie wavelengths are not relativistically covariant as simultaneous spatial structures, making any such wave properties physically inconsistent. In this work, we explore whether modern experiments vindicate an alternative view: that apparent waviness in diffraction and interference scenarios emerges as a consequence of quantised interactions between particles. Such a view has historically received very little attention, despite being the exact modern explanation of both the Kapitza-Dirac effect and ultrafast electron diffraction. We then study a photon orbital angular momentum realisation of the double slit to prove this explanation capable of unifying barrier and phase-grating diffraction phenomena. Finally, we demonstrate that the quantum formalism demands that particle momentum is determined at the point of scattering, contravening wave accounts of quantum interference.
1 Introduction
Feynman is often quoted as describing the double slit as containing “the only mystery” in quantum mechanics [1]. This experiment is indeed curious, apparently requiring that otherwise material particles behave as waves some of the time. Such results inspired Bohr’s complementarity: that neither particle nor wave is sufficient to encapsulate quantum entities alone. However, there are good reasons not to be sanguine about the physical reality of this waviness. First, we can note that these effects are apparent in the statistics of quantum experiments, with individual runs showing particle properties instead. As Popper has noted, in assigning wave nature to individual particles via their bulk statistics we commit the ecological fallacy [2]. This leads to the clumsy resolution of “wavefunction collapse”, semi-dead cats [3], and the interminable measurement problem [4, 5]. A further concern might be application of the Huygens-Fresnel principle to probability waves that cannot be disturbances in any sense.
There is, however, a far graver problem with quantum waviness: the wavelengths are not covariant as correlation lengths in any form of relativity. In 1975 Landé argued [6] that because de Broglie wavelengths take the form $\lambda=\frac{h}{p}$ they must vary with inertial frame. Landé also assumed that all real waves should possess phase invariance, generating an apparent paradox. Lévy-leblond [7] showed that $\lambda$ varying with inertial frame is necessary to maintain the Galilean invariance of Schrödinger equation solutions and thus phase invariance does not hold for quantum waves. This establishes that $\lambda=\frac{h}{p}$ is not a Galilean invariant, indeed it is manifestly not covariant as a length in special relativity either. By this we mean that we cannot measure $\lambda$ as the simultaneous length of an extended structure in space. This is not the case for sound or water waves. So $\lambda$ cannot be interpreted as a simultaneous physical length scale. The solution in relativistic quantum theories is to rely on $p^{\mu}=\hbar k^{\mu}$ as a definition of a covariant wave-vector. This appears to leave us with a paradox. We can provide a covariant mathematical description, but the wavelength is not covariant in the manner we should expect. We should clearly ask what kind of covariance does $\lambda$ then exhibit. In answer, we consider the Lorentz transform for distances propagated by an object with speed $u\to c$, in some reference frame $S$, over a time $\Delta t$ or
$\Delta x^{\prime}=\Delta x\sqrt{\frac{1-\beta}{1+\beta}}\;,$ (1)
where $\beta$ is the speed of frame $S^{\prime}$ relative to $S$. The transformation has the same structure as the relativistic Doppler shift for electromagnetic wavelengths, providing an indication that such wavelengths transform like photon propagation distances, not the correlation length of waves extended in space. Moreover, a Doppler-like factor occurs similarly for any propagation speed $u$: $\Delta x^{\prime}=\gamma\Delta x\frac{(u-v)}{u}$. Further, we can express $p^{\prime}=\gamma p\frac{u-v}{u}$ for particle moving with $u\leq c$, which yields $\lambda^{\prime}=\frac{\lambda}{\gamma}\frac{u}{u-v}$, with the Galilean form being $\lambda^{\prime}=\frac{h}{m(u-v)}=\lambda\frac{u}{u-v}$. Repeating our pattern of similar structure to the transformation of propagation distances. Notably, our wave-particle dilemma only occurs for cases where the wavelength transforms like a propagation distance. Material waves (e.g. water or sound) do not carry this Doppler factor on $\lambda$, merely experiencing length contraction during Lorentz boosting.
We now have a resolution of our “paradox”. The de Broglie wavelength appears non-covariant if we wish it to be an actual length of a structure that exists simultaneously in space (like a wave). However, if we regard it as a propagation distance instead, with no inherent simultaneity in any frame, there is perfect accord with relativistic covariance. Thus, in using $\lambda$ as a length-scale to explain phenomena like interference and diffraction of quantum objects we are tacitly violating relativistic covariance. It is therefore not possible to ascribe actual waviness to quantum objects in a physically consistent manner, as is similarly concluded in [8] (although our argument here is different). All of this should cast some doubt on the physical, rather than mathematical, validity of de Broglie-Bohm theories, and potentially also on the correspondence between field modes and particles used in quantum field theory. A manifestation of the former may be seen in [9], where real pilot waves fail to reproduce the double slit experiment, possibly implying only those with physically inconsistent wavelengths can do so.
Further evidence that real waviness is unnecessary can be found in experiments where the double slit is performed in single atoms [10]. Here rubidium is photo-ionised by independent lasers of different frequency that, in combination, offer two possible ionisation paths. The resulting interference of photo-electron waves is understood to occur only in the abstract state-space, making it difficult to conclude that this interference requires actual waviness. The authors note that it is the quantised atomic transitions that provide phase coherence in the first place. This provides motivation to explore the role of quantisation as an explanation for phenomena that otherwise look like interference patterns formed by extant waves.
In particular, this work highlights a long-overlooked means to understanding apparent waviness in matter: quantised interactions. If incoming particles can only exchange discrete quanta of momentum with periodic scatterers (like slits or diffraction gratings), then we can describe “interference” patterns by local particle interactions following Duane’s 1923 arguments [11]. Here the grating is a quantum system, with allowed momentum states quantised according to its periodicity. Incoming particles interact with it and acquire quantised momentum along the same axis as the periodicity. Trajectories whose momenta match the quantisation rule arrive at “constructive interference” regions, while those requiring forbidden half-quanta yield “destructive interference”. In fact, a very similar approach is used for explanation of the Kapitza-Dirac phase-grating effect [12]. Here electrons are diffracted by a standing wave laser field. The diffraction peaks are sourced from absorption of quanta from the standing wave.
Indeed, Landé proposed Duane’s hypothesis as a solution to the de Broglie wavelength’s lack of wave-like covariance [13, 6]. His arguments, however, amount merely to illustrating mathematical consistency, along with the allegation that his thesis is no more absurd than the quantum orthodoxy [6]. Once this is combined with the objection that local interactions cannot allow particles to “experience” an extended structure, like a diffraction grating, his arguments hardly seem compelling. This work aims to bring Duane’s hypothesis into a new century with compelling evidence that was not available to earlier proponents. This demonstrates that serious, realist interpretations of quantum phenomena are possible, and require no extension of, nor additions to, the formalism. In this sense, we present an argument similar to that in [14], but without association to Bohmian extensions of quantum mechanics.
We note some similarity to the double-slit explanation proffered in [15], where particles traverse only a single slit and do not “self-interfere”. The difference can be understood in terms of the usual question: “how does the particle know about a slit it doesn’t traverse”. Conventionally, waves are the answer. In [15], it is non-local particle properties. However, our picture is that the slits’ quantum structure places limits on the allowed local interactions with incoming particles. Thus, a two-slit structure’s difference to a single slit requires no non-local effects, unless all phenomena described by phonons in condensed matter are also deemed non-local. Additionally, it has been argued that diffraction of delocalised matter waves, on the molecular scale, requires accounting for local interactions between the molecule and diffraction element [16]. We extend this work by arguing it is true even for point particles.
This paper is structured as follows: in section 2 we explore the explanation of the Kapitza-Dirac effect and how it has an elegant description in terms of local particle interactions. Then, in section 3 we explore how this description can be extended to other scenarios where diffraction and interference occur. To prove that quantised particle interactions are a viable alternative to interference we explore a double-slit experiment with photon orbital angular momentum in section 5. Then we demonstrate that quantum mechanics requires the far-field double slit pattern is determined at the slits in section 6 and discuss coherence as well as “which way” information in section 7. Finally, we draw conclusions in section 8.
2 The Kapitza-Dirac effect
This phenomenon, proposed in 1933 [17] and later confirmed [18, 19], is where an electromagnetic standing wave acts as a diffraction grating for charged (or polarisable) particles. Ostensibly this represents a symmetry between light and matter, which is a reflection of the particle-wave duality. However, the KDE has a wave-particle object diffracted by a wave, whereas conventional diffraction has a wave diffracted by a classical barrier. This seems decidedly non-dual. The standard explanation is we are discussing a phase grating, i.e. spatially varying phase differences are applied to the matter wave by the standing light wave due to the electromagnetic coupling of the particles [12]. We can immediately see that this is not actually explained by waves, both wave and particle properties are needed.
However, the presence of diffraction maxima in the KDE can be fully explained in terms of incident particles absorbing discrete light quanta of fixed energy [12]. In particular, the particle (travelling initially in the $x$-direction) absorbs a photon with $y$-direction momentum $p_{y}=\hbar k$ ($k$ is the laser wavenumber) and subsequently emits a photon at $p_{y}=-\hbar k$ via stimulated emission from the reflected laser beam. Thus, the incident particle incurs a momentum change of $2\hbar k$. Since this can happen multiple times as the particle crosses the interaction region, we would expect a momentum spectrum for the out-going particles to be given in terms of $2n\hbar k$, where $n$ is an integer. The intensity peaks are indeed found to correspond to these values.
We can explain the KDE completely in terms of particle interactions where the diffraction element is an active participant, we do not need to invoke any waviness. This is achieved by treating the diffraction element itself as a quantum system. Notably, we cannot actually describe the KDE without reference to particle properties, as even a wave picture must use particle interactions with the electric field to explain the wave’s phase shifts [12]. However, in barrier diffraction, a wave explanation treats the grating as a completely classical barrier to propagation and requires no reference to particle properties except at the point of detection. Perhaps we can symmetrise the KDE and other diffraction scenarios with a quantum description of the diffraction element? If possible, we would no longer require the physically inconsistent wave explanation. Furthermore, it would show that assigning a wave nature to quantum objects is contingent on regarding the diffraction element as classical.
3 Diffraction by material gratings
We turn then to more conventional diffraction phenomena, electromagnetic or matter waves diffracted by material gratings. These scenarios apparently demand a wave-like description, even if this requires that a particle’s wave aspect “self-interferes” and suffers from our subtle inconsistency with all forms of relativity. To facilitate our search for a picture where the diffraction element is no longer a passive, classical barrier, we will consider a 1D diffraction grating, centred at $x=0$ and $y=0$, that extends symmetrically in the $y$ direction. A distant screen is used to detect diffracted particles at $x=L$, where $L\gg d$ (the grating periodicity scale). It is well known that the diffraction maxima of a periodic grating can be found at angles $\theta$ (measured from the $x$-axis) given by
$d\sin\theta=n\lambda\;,$ (2)
where $\lambda$ is the de Broglie wavelength, and $n\in\mathbb{Z}$. This can re-imagined via the de Broglie relation as
$p\sin\theta=\frac{nh}{d}\;.$ (3)
Recognising that $p\sin\theta=p_{y}$, we can see that this is a quantisation rule on the transverse momentum of the outgoing particle. Particularly, interaction between incident particle and grating must involve the quantised exchange of momentum. Simply by reinterpreting a wave-interference formula, we have found an explicit discussion of the grating as participating in a quantum interaction. Our use of the de Broglie formula is merely heuristic, and we will see in section 6 that the momentum expression is the fundamental one extracted from the quantum formalism. Our interacting picture does not merely explain the diffraction maxima, as the minima can be similarly identified by
$p\sin\theta=\frac{mh}{2d}\;,$ (4)
where $m$ is an odd integer. Thus, the forbidden scattering angles, that apparently required a wave-based explanation, can be understood as being ruled out by requiring a forbidden half-quantum exchange of momentum. To elaborate, the $y$-momentum needed for deflection of the particle to “destructive interference” positions cannot be obtained from interactions with the periodic scatterer or its constituents, due to quantisation of collective momentum states (to be justified below). This quantisation rule was first observed by Duane [11] in 1923, but has only received scattered mentions in the literature over the last century [20, 13, 21, 6, 4, 22, 23, 24]. The generalised version of the simple analysis above fully reproduces both Laue and Ewald conditions [21]. Notably, this viewpoint requires that the diffraction grating must be characterised as inherently quantum. However, this comes with a difficulty: how to justify the quantised nature of the interaction? Mathematical justification is easy to find [22, 23]: Bloch’s theorem [25] requires that structures characterised by spatially periodic potentials have quantised momentum eigenstates. For a crystalline solid, these are collective motion states or phonons. The scattering interaction could therefore be similar to the physics surrounding the Mössbauer effect [26], where the incident particle interacts with a local lattice site and recoil can be absorbed by collective motion [27, 28]. The interactions that mediate collective motions naturally restrict local recoils as well, hence the points of “destructive interference” corresponding to forbidden lattice momenta. Thus, the intensity peaks correspond to resonant scattering, where exact lattice quanta are exchanged, and the continuity is supplied by local recoils. Notably, the explanation of the Mössbauer effect eliminates the problem that stymied Landé [6]: it is possible for a local interaction to depend upon the global structure of a participant. Importantly, this explanation hinges upon the periodic structure of the diffraction grating, with Bloch’s theorem assuring us it will extend to any periodic diffraction phenomenon and [22] extending it to less symmetric structures. However, there is no reason that the same physical mechanism should explain every example of such interference. Symmetry guarantees interference mathematically, but does not uniquely specify any physics.
The looming question is whether such a model can be compatible with the uncertainty principle, as it is commonly held that a particle cannot possess definite momentum and position at the same time (see e.g. [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 1]). This is not a problem for Duane’s model, as the aforementioned belief is a fallacious use of the principle [4, 2] with many more conventional texts acknowledging this [40, 41, 42, 43, 44]. As rigorously defined, the uncertainty principle applies to standard deviations over a measurement ensemble [45], it thus has nothing to say about the ontic status of individual particle properties. Indeed, Popper argued that were the principle to apply to individual particles, the minimum ensemble deviation could never be probed [2].
4 Evidence supporting Duane’s model
Interestingly, there are experiments that already exhibit a hallmark of Duane’s mechanism: the sensitivity of diffraction patterns to excitation of phonon modes within the diffraction element prior to scattering [46, 47, 48]. The explanation of these results is explicitly done in terms of particle-phonon scattering interactions but is often written in “wave language”. For instance, in [46, 47], the authors describe the exchanged wave number $Q$ by $Q=q+K_{Q}$, where $q$ is the phonon wave number and $K_{Q}$ is the closest reciprocal lattice vector to $Q$. Thus, when no phonons are excited, the exchange is quantised according to the reciprocal lattice vector, which follows from the general form of Eq. (4) [21, 22]. When photons are excited, the situation is very closely analogous to the KDE, we have an incident particle diffracted by a phonon wave, and we cannot explain this by resorting to wave-wave diffraction without referring to particle interactions. However, we have the option of a complete, physically consistent, description in terms of local particle-particle interactions. Here the phonon is understood as a representation of quantised states of collective motion enforced by the lattice symmetries, interaction with such states must then be mediated via local interactions at lattice sites. Thus, advanced diffraction experiments with practical applications already use Duane’s picture without even realising it.
A further point of interest arises in [47]. Here, it is demonstrated that exciting phonons alters the diffraction spectrum in a very particular manner. The apparent effect is that phonon excitation broadens the Bragg peaks normally associated with the target crystal. The more rapidly the probe follows the excitation, the more pronounced this effect is. What this demonstrates is a case where continuity is supplied between otherwise discrete peaks. This is important, as it demonstrates an analogy to precisely the question we should ask about Duane’s mechanism: why is the spectrum continuous if the momentum states are strictly quantised? While it is unlikely to be the same mechanism that answers the question in general, it does suggest that physical processes exist that perform the required role.
In addition, it has also been shown that van der Waal’s interactions during diffraction contribute to the pattern, in a manner degenerate with the slit width [49] (similar phenomena are used to argue against wave-particle duality in [16]). In these experiments, beams of molecules are diffracted by silicon nitride gratings. The diffraction patterns are found to depend on the dipole polarisability of the incident molecules. As previously mentioned, the effect of the extra interactions produces a modified “effective slit width”. This is very suggestive of Duane’s explanation, which connects slit width to interactions already. Interestingly, these additional interactions also lift the zeroes of the diffraction pattern. This is unsurprising if diffraction momentum changes are being supplied by interactions in general, as the van der Waal interactions can allow deflection angles otherwise forbidden by the intrinsic lattice momenta. On the other hand, it is quite mystifying if this is actually wave interference at work. The wave boundary conditions are not being altered, yet the interference patterns change. More crucially, this case supplies direct evidence that local interactions within diffraction gratings can supply an effect that is degenerate with the boundary conditions of apparent wave interference, precisely what is being claimed by Duane.
It is evident that we can discuss all of our diffraction scenarios in terms of particle interactions. However, we are now in a situation where the necessity of particles having wave properties is contingent on classical descriptions of the diffraction element. Additionally, there are scenarios where this description cannot avoid inclusion of particle interactions. If we consider the fully quantum description more fundamental, this reduces the wave explanation to an emergent consequence of quantisation, with only heuristic value. Our next step will be to introduce further evidence that only a particle-particle interaction description is physically consistent.
5 The angular double slit
Duane’s picture is not limited to diffraction gratings. Indeed, it was quickly shown to apply to the double slit as well [20]. Modern experiments can offer us an analogy to demonstrate that quantised exchanges can produce “interference patterns”. To do this, we consider a double slit experiment where the momentum analogue is inherently quantised. Thus, it can only be changed by discrete quanta. If we can obtain the analogous phenomena to a regular double slit, we have unambiguous, empirical evidence of the viability of Duane’s mechanism.
To this end, we can construct an analogue to the double slit using the Orbital Angular Momentum (OAM) of light, $\ell$, as is done in [50]. In the angular double slit, a spatial light modulator displays two angular sections of a fork hologram. The net result is the incident beam receives OAM from both the angular slits and the fork hologram (the latter allowing for spectral decomposition of the outgoing beam). The spectral prediction is given by [50]
$P(\ell)\propto\mathrm{sinc}^{2}\left[\frac{\beta\ell}{2}\right]\cos^{2}\left[\frac{\alpha\ell}{2}\right]\;,$ (5)
where $\alpha$ is the slit spacing and $\beta$ their width. This is verified empirically [50] and can be compared to the linear form
$P(p_{y})\propto\mathrm{sinc}^{2}\left[\frac{\sigma p_{y}}{2\hbar}\right]\cos^{2}\left[\frac{p_{y}\Delta y}{2\hbar}\right]\;,$ (6)
where $p_{y}$ is the transverse momentum, $\sigma$ is the slit width, and $\Delta y$ is the spacing. Writing $\alpha=f\pi$, a simple correspondence can be established between the two by introducing the unit-free momentum $\tilde{p}=2\frac{p_{y}\Delta y}{fh}$ so that
$P(\tilde{p})\propto\mathrm{sinc}^{2}\left[\frac{f\pi\tilde{p}}{2}\frac{\sigma}{\Delta y}\right]\cos^{2}\left[\frac{f\pi\tilde{p}}{2}\right]\;.$ (7)
The factor of 2 in $\tilde{p}$ corresponds to the 2-fold rotational symmetry of the angular mask [50].
A vital aspect of this experiment is that $\ell$ modes are “free-space modes” [51]. Meaning they do not change $\ell$ while propagating in free space. Thus, a beam’s angular momentum is set by a discrete exchange at the spatial light modulator and not determined at some point of overlap at the detector. This is consistent with the ability to perform the experiment with classical beams. Thus, our $\ell$ spectrum can be simply interpreted as the fraction of beam power propagating in mode $\ell$.
We display an example OAM spectrum from such an arrangement in Figure 1. This pattern can be simply understood as resulting from an interaction between the slits and beam that imparts discrete units of $\ell$. It is notable that our $\ell$ spectrum, although discrete, is perfectly matched by the linear double slit far-field intensity pattern (or momentum representation distribution). This analogy provides evidence that a discrete exchange interaction can, in empirical fact, produce something that looks like an interference pattern. Since $p$ is not inherently quantised, but reproduces the case where $\ell$ is, we have to conclude that quantisation is being imposed upon $p_{y}$ via interaction with the slits.
[…]
Figure 1: OAM spectrum from angular double slit when $\alpha=\pi$, $\beta=\frac{\pi}{3}$ plotted as a bar graph. The line plot shows a linear double-slit momentum distribution in $\tilde{p}$ when $\frac{\Delta y}{\sigma}=3$ and $f=1$.
6 The double slit revisited
An intriguing aspect of the OAM double slit in section 5 is that it does not matter (in principle) where we measure the $\ell$ spectrum. Both near-field and far-field diffraction regimes yield the same spectrum. For wave-based explanations this result is extraordinary. The “interference” pattern we see in the far-field is already imprinted in $\ell$ (or by analogy $p_{y}$) within the near-field. It is easy to demonstrate that this is so. Consider a position wavefunction at the double slit position $\psi(\vec{x})$. This can be written in the momentum representation in terms of Fourier coefficients $a_{k}$ (for each momentum $k$). The momentum spectrum is then $|a_{k}|^{2}$. In this same plane wave basis, propagation in free space merely accumulates a phase factor, or $\psi(\vec{x}+\vec{\delta x})=\int a_{k}\mathrm{e}^{i\vec{k}\cdot\vec{x}}\mathrm{e}^{i\vec{k}\cdot\vec{\delta x}}dk$ after propagating $\vec{\delta x}$. Thus, the Fourier coefficients, and hence the $p$ spectrum, are unchanged by free-space propagation. The same argument holds for the angular case when using, for instance, the Laguerre-Gauss basis [51]. The fact that changes occur in the position space pattern, but not the momentum spectrum is then down to each mode accumulating differing phases. We will revisit this from a particle point of view later.
What the above argument demonstrates is that a consistent momentum distribution undergirds interference patterns, rather than requiring the pattern be determined by wave-overlap at the point of detection. To make this complete, we can demonstrate that the quantum formalism itself requires that the momentum spectrum is determined before wave over-lap can even occur.
We start by formulating a simple, position-representation amplitude for particles passing in the $x$-direction through the linear slit apparatus (positioned at $x=0$ and extending symmetrically in the $y$-direction)
$\psi(y)\propto\Pi\left(\frac{y-\Delta y/2}{\sigma}\right)+\Pi\left(\frac{y+\Delta y/2}{\sigma}\right)$ (8)
where $\Pi\left(\frac{y}{a}\right)$ is the rectangular function of width $a$, centred at $y=0$. It should be noted that there are no interference terms contributing to $|\psi|^{2}$, which just corresponds to an unremarkable distribution for particles occupying any point within either of the two slits with equal probability. This $\psi(y)$ is valid only at the slits’ position ($x=0$) on the propagation axis. The momentum amplitude is then found via the usual Fourier transform
$\phi(p_{y})\propto\int_{-\infty}^{\infty}\left[\Pi\left(\frac{y-\Delta y/2}{\sigma}\right)+\Pi\left(\frac{y+\Delta y/2}{\sigma}\right)\right]\mathrm{e}^{\left(-i\frac{\vec{x}\cdot\vec{p}}{\hbar}\right)}dy\;.$ (9)
This results in
$\phi(p_{y})\propto\mathrm{sinc}\left[\frac{\sigma p_{y}}{2\hbar}\right]\left(2+\mathrm{e}^{i\frac{\Delta yp_{y}}{\hbar}}+\mathrm{e}^{-i\frac{\Delta yp_{y}}{\hbar}}\right)\;,$ (10)
where any contributions from orthogonal momenta amount merely to global phase factors. The Fourier integration over all $y$ can be understood here as considering the global structure of the scatterer, rather than invoking interfering plane waves. Our momentum distribution is then
$|\phi(p_{y})|^{2}\propto\mathrm{sinc}^{2}\left[\frac{\sigma p_{y}}{2\hbar}\right]\cos^{2}\left[\frac{p_{y}\Delta y}{2\hbar}\right]\;.$ (11)
This indicates that the $p_{y}$ spectrum, with both single and double-slit factors, is already imposed at $x=0$, before waves originating from each slit get a chance to overlap. Thus, actual spatial wave-overlap and interference is irrelevant in forming the far-field position distribution, which precisely mirrors $P(p_{y})$ above. This conforms nicely with the $\ell$ analogy, as the beams in question do not change $\ell$ while propagating in free space (note the experiment can be done entirely classically). Meaning their $\ell$ must only be determined at the spatial light modulator, not in some interference region at the detection point (unless one wishes to argue the detector itself couples different modes based on its distance to the slits). Moreover, the maxima of $\cos^{2}\left[\frac{p_{y}\Delta y}{2\hbar}\right]$ occur when $p_{y}=\frac{nh}{\Delta y}$, the same quantisation rule observed by Duane, as $\Delta y$ is the symmetry scale of the slits. Although this is a reflection symmetry, rather than a full periodicity, momentum quantisation appears following similar arguments [22]. Additionally, the factor $\mathrm{sinc}^{2}\left[\frac{\sigma p_{y}}{2\hbar}\right]$ has maxima where $p_{y}=\frac{mh}{2\sigma}$ where $m$ is an odd integer. These maxima are spaced out by integer values of $\frac{h}{\sigma}$ and the minima occur $\frac{h}{2\sigma}$ from the maxima. Since $\sigma$ represents a second symmetry scale, the presence of two simultaneous quantisation rules is unsurprising. The “missing orders” occur when minima $\frac{m^{\prime}h}{\sigma}$, $m^{\prime}\in\mathbb{Z}$ match maxima $\frac{nh}{\Delta y}$. If we let $\sigma=f\Delta y$ this matching condition means when $\frac{m^{\prime}}{f}$ is an integer we find a “missing order” that was disallowed by the $\sigma$ scale quantisation.
What then is the particle equivalent of the mode-dependent phase accumulated during propagation? For waves, this phase determines the variation of the position-space image with distance between the slits and screen. The obvious answer is to consider whether the particle’s position in the slits affects its destination. This would be apparent in the near-field, but averaged out in the far. Thus, we could simply make geometric projections for particles originating from different points in the slits, each sampled from the same momentum distribution above. However, we will not reclaim the full detail of a Fresnel pattern. This is because the derivation above does not expose the details of interactions between incident particles and slits [24]. In that, there is no explicit in-slit position dependence. We should expect particles passing closer to the edges have differing interaction probability with each edge and will contribute to the spectrum differently to a particle passing exactly between both edges. Thus, our above result is the correct total spectrum (integrated over the slit width), but lacks the detail that would result in something of similar complexity to a Fresnel integral for the near-field diffraction pattern. The different propagation phase for each wave mode is thus tracing the in-slit position association of each mode. This also provides a very natural explanation of why diffraction images converge to a far-field form, as this is the long-distance geometric projection of trajectories sampled from the momentum spectrum.
It is clear then that the interference-like behaviour in the diffraction of quantum systems can be consistently viewed as originating from the quantisation of particle momenta. This is evidenced directly by the angular double slit, as well as other experiments invoked in section 4. Additionally, the momentum spectrum is determined at the slits and then remains unchanged during propagation (as shown by the far-field pattern reproducing it) in complete harmony with Duane’s hypothesis. This holds even when considered within the mathematics usually attributed to the wave nature of quantum objects. Notably, there is no amplitude interference in the position picture, but it is present in that of momentum. The Fourier transform is viewed here as an assessment of the global periodic structure of the scatterer, rather than a collection of interfering plane waves. The status of momentum can be further reinforced by its conservation for individual recoil events in double slit systems [52]. Such an effect could only occur retrocausally in any interpretation where the interaction occurs via waves. That is, after the particle’s momentum is determined by capture at the screen, but is perfectly natural in Duane’s picture.
7 Coherence and “which way” information
The double slit experiment requires light which exhibits spatial coherence, it is natural to ask how this requirement is interpreted without invoking waves. The coherence between two waves or fields is commonly described as their statistical similarity [53]. This is reinforced for quantum coherence as there is some correspondence between it and entanglement [54, 55]. This has a very natural interpretation in the language of particles: we need a particle distribution whose momentum statistics will not obscure the effect of the double-slit scattering itself. More generally, a multimode photon source has a coherence length because differences in photon momenta become apparent over a long enough distance, hence the dependence of a coherence length on bandwidth and the speed of light. In a diffraction scenario this would produce blurring via incompletely overlapping patterns for each mode.
Another interesting case is if we separate the double slits far enough apart, the cosine squared in Eq. (11) oscillates rapidly between $1$ and 0 with the overall pattern being two single slit diffraction results. This is sometimes discussed as an effect of coherence. However, we see it is because $\frac{h}{\Delta y}\ll p_{y}$, i.e. it is a continuum limit for the double-slit quantisation where the fringe spacing goes to 0 compared to the single slits. The comparison to $p_{y}$ explains why the wavelength is often compared to $\Delta y$ to determine an appropriate slit separation to see double-slit interference, as we can reformulate our condition to $\frac{h}{\Delta y}\ll\frac{h\sin\theta}{\lambda}$ or $\Delta y\gg\lambda$ (as $\sin\theta$ is bounded).
A final point of interest is “which way” information. This is a phenomenon where determining which slit a scattered particle traversed can remove the interference pattern [1]. More interestingly, an uncertain determination of the path merely reduces the contrast of the pattern in proportion to the uncertainty [56]. This appears to be strong argument for wave behaviour, as by ‘marking’ the waves emanating from different slits, we are removing the essential coherence needed to produce interference. However, it is very natural in particle language as well. To illustrate, we will consider the case of determining “which path” information by scattering photons off of particles as they emerge from the slits. We choose this case for physical transparency. It is well known that the spatial resolution of such scattering detections is limited by probe wavelength. However, this can be easily proven for a pure particle approach. All we need do is regard a particle emerging form the slits and absorbing a probe photon as being excited. This excitation has an average lifetime $\tau$ that obeys $E\cdot\tau\geq\frac{\hbar}{2}$ where $E$ is the excitation energy. Thus, the position uncertainty of this probe is the distance moved as a result of absorption of the probe photon prior to re-emission, or $\Delta x=\Delta v\tau$. Here $\Delta v$ is the velocity kick introduced by absorption. Now we can write
$\Delta x\geq\Delta v\frac{\hbar}{2E}=\frac{\Delta v\hbar}{m\Delta v^{2}}\;.$ (12)
Where the second equality assumed that we can ignore the initial trajectory of the particle (we are only interested in the excitation and can happily choose the initial rest-frame of the particle). Since $m\Delta v=p_{\gamma}$, or the probe momentum, we have that $\Delta x\geq\frac{\hbar}{p_{\gamma}}$. Now we can see that discriminating between slits will require photons with momenta $\sim\frac{h}{\Delta y}$, very similar to the momentum quantisation responsible for the apparent interference pattern. Moreover, we can imagine that the particle’s $p_{y}$ is randomly changed during the probe scattering, as there are no preferential scattering angles. We can implement this by convolving $P(p_{y})$ with a flat window-function between $p_{y}-\frac{h}{\alpha\Delta y}$ and $p_{y}+\frac{h}{\alpha\Delta y}$ where $\alpha\in\mathbb{R}$. Higher resolution, i.e. $\alpha<1$, results in a broader convolution that reduces the contrast of the bright and dark fringes.
8 Conclusion
The de Broglie wavelength is not covariant as the length scale of a simultaneous extended structure in any reference frame. Thus, we cannot use it as physical explanation of observed phenomena, it is merely a useful heuristic. This puts claims about wave-particle duality under extreme strain and undermines otherwise compelling explanations of interference and diffraction. To resolve this, we have argued that the apparent wave nature of quantum objects is only manifest when diffraction elements are treated classically. In contrast to the Kapitza-Dirac effect, which never needs to invoke either the waviness of incident electrons or the scattering photons. When the diffraction element is considered as a participating quantum system, a picture of diffraction and interference emerges that describes them as apparent consequences of quantised momentum exchanges. Thus, uniting phase-grating and barrier diffraction experiments with one physically consistent explanation. This is justified by the periodic structure of the scatterer implying it must have quantised momentum eigenstates. Moreover, a double slit analogue with the explicitly quantised photon orbital angular momentum shows the same expected distribution as the linear double slit momentum. This provides evidence that a quantised exchange interaction can produce what looks like an interference pattern. Moreover, the only conclusion to be drawn from the analogy is that the interaction is responsible for quantising linear momentum in a double slit. Vitally, it is shown that the momentum distribution is determined at the slits themselves, not at the point of particle detection. Finally, we also presented evidence that this viewpoint is already used implicitly by experimenters making use of electron diffraction to characterise phonons in condensed matter and this is the only physically consistent way to view these experiments.
This work demonstrates that complex amplitude interference is capable of representing particle physics that it does not resemble at all. Thus, explaining why naive positivism resulted in a view of quantum mechanics at odds with realism and logic. Moreover, we see an indication of a relationship between quantisation and apparent wave behaviour being mediated via complex amplitude interference, meaning that the heuristic value of waviness is not just a deceptive coincidence.
It is not immediately clear how this viewpoint generalises to interference phenomena related to beam splitters. Such a study would hinge upon a description of beam-splitters as active participants in interactions with photons, something beyond the scope of this work. However, our conclusions here raise the possibility that a similar realist interpretation could be buried beneath the obscuring mathematics.
Acknowledgements
Thanks to Robert de Mello Koch, Sam van Leuven, and Andrew Forbes for helpful discussions; as well as Emil Roduner and Tjaart Kruger for their encouragement.